OpenStudy (anonymous):

Sara selects two of the seven colors of the rainbow, one after another, with replacement. What is the probability that the first color is red and the second is orange?

OpenStudy (carolina4567):

1 /49

OpenStudy (carolina4567):

see prob of selecting red is 1/7 and after replacement probability of selecting orange is again 1/7 hence reqd =1/7 * 1/7 =1/49